### v2

# A universal black-box quantum Monte Carlo approach to quantum phase transitions

Abstract
We derive exact, universal, closed-form quantum Monte Carlo estimators for finite-temperature energy susceptibility and fidelity susceptibility, applicable to essentially arbitrary Hamiltonians. Combined with recent advancements in Monte Carlo, our approach enables a black-box framework for studying quantum phase transitions — without requiring prior knowledge of an order parameter or the manual design of model-specific ergodic quantum Monte Carlo update rules. We demonstrate the utility of our method by applying a single implementation to the transverse-field Ising model, the XXZ model, and an ensemble of models related by random unitaries.
I Introduction
Formally, a quantum phase transition (QPT) is defined by nonanalytic behavior of a system’s ground-state energy in the thermodynamic limit [1]. Traditionally, QPTs are analyzed using local order parameters, such the average local magnetization in the transverse-field Ising model (TFIM) [1]. However, identifying a suitable local order parameter for a new model is often challenging—and for topological QPTs, it is entirely infeasible [2, 3]. Consequently, substantial effort has been devoted to developing generic methods for studying QPTs that do not rely on prior knowledge of an order parameter [4, 5, 6, 7, 8, 9, 10, 11, 12, 11, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Nevertheless, these approaches still rely on manually encoding subtle, system-specific prior knowledge. Building on recent advancements in quantum Monte Carlo (QMC) methods [23, 24, 25, 26] and the well-established fidelity-based approach to QPTs [4, 17, 18, 22, 19], we propose an automated, order-parameter-free, black-box method for studying QPTs that requires no manual encoding of system-specific details.
More precisely, we derive QMC estimators for energy susceptibility (ES) [18, 27, 28] and fidelity susceptibility (FS) [4, 9, 17, 18, 19], which can identify quantum critical points in QPTs without relying on order parameters. While related estimators have appeared in other QMC frameworks [18, 17, 19], our derivations and resulting estimators are developed within the relatively recent permutation matrix representation QMC (PMR-QMC), which offers several key advantages over earlier methods [29, 30, 31, 23, 24, 25, 26]. Most relevant to this work, it was recently shown that ergodic PMR-QMC update rules satisfying detailed balance can be generated via a deterministic, automated algorithm for essentially arbitrary Hamiltonians [23, 24, 25]. Furthermore, our estimators are applicable to arbitrarily complex driving terms, as we demonstrate in this work. To the best of our knowledge, neither of these capabilities was supported by previous approaches. Taken together, these features allow our method to be applied directly to a wide range of systems without the need to manually encode laborious, system-specific details.
To this end, we apply our approach to estimate ES and FS in three spin systems: a 2D TFIM model, a 2D XXZ model, and a 100-spin model with random Pauli operator terms that—to our knowledge—cannot be studied using any other method. Given any spin-$1/2$ Hamiltonian expressed as a sum of Pauli operator products, our current implementation [32] can attempt to locate quantum critical points by estimating ES and/or FS. As with any QMC method, our approach may encounter convergence issues due to frustration or the sign problem [33, 34], which can sometimes be mitigated with custom QMC updates [35, 36] or other techniques [37, 38, 33, 39, 40, 41, 34]. For the three models studied in this work, however, no manual, system-specific adjustments were required to ensure convergence. In all cases, our code automatically monitors the severity of the sign problem [42, 23] and performs thermalization and autocorrelation diagnostics to ensure reliable estimates.
Our implementation [32] builds upon preexisting open-source code for arbitrary spin-$1/2$ models [42]. Since our estimators are derived within the general PMR-QMC framework, which can represent arbitrary Hamiltonians [29, 26], they can readily be ported to any existing or future PMR-QMC codebases. At present, this includes Bose-Hubbard models on arbitrary graphs [24], arbitrary spin-$1/2$, high-, and mixed-spin systems [23, 25], and fermionic systems [25, 43]. In summary, PMR-QMC can automatically generate QMC moves for essentially any Hamiltonian [23, 24, 25], and in this work, we derive generic ES and FS estimators within this framework. Combined with automatic convergence diagnostics, our approach—unlike previous numerical methods—can reliably be used as a black-box tool for studying QPTs.
II Energy and fidelity susceptibilities
We discuss the ES and the FS as two generic, order-parameter-free tools to study QPTs [18, 19, 4, 9]. Throughout, we assume an $n$ spin Hamiltonian with non-degenerate groundstate,
$H(\lambda)=H_{0}+\lambda H_{1},$ (1)
that has quantum critical point (QCP) at $\lambda=\lambda_{c},$ denoting the boundary between two quantum phases of a QPT. Henceforth, we refer to $H_{0}$ as the static term and $H_{1}$ as the driving term for convenience. Given the spectral decomposition, $H(\lambda)\ket{\psi_{k}(\lambda)}=E_{k}(\lambda)\ket{\psi_{k}(\lambda)},$ we define the ES [18, 27, 28]
$\chi_{E}(\lambda)=-\frac{\partial^{2}E_{0}(\lambda)}{\partial\lambda^{2}}.$ (2)
For second-order QPTs such as that in the TFIM [1], the ES is expected to have a peak or even diverge, and hence, it can be used to locate QCPs. Similarly, the second-order term in the Taylor expansion of groundstate fidelity, or FS [9],
$\chi_{F}(\lambda)=\lim_{\delta\lambda\rightarrow 0}\frac{-2\ln|\langle\psi_{0}(\lambda)|\psi_{0}(\lambda+\epsilon)\rangle|}{\delta\lambda^{2}},$ (3)
also exhibits a peak or diverges at QCPs for a wide variety of QPTs [28, 44, 18], including first-order [11], second-order [11], topological [45, 46], and as seen recently, even dynamical QPTs [11].
There are several equivalent characterizations of both quantities that are useful in different contexts. First, we discuss various rich characterizations of the FS [4, 9, 44, 28, 7]. By a direct Taylor expansion and simplification of $\langle\psi_{0}(\lambda)|\psi_{0}(\lambda+\delta\lambda)\rangle,$ one finds the differential form,
$\chi_{F}(\lambda)=\langle\partial_{\lambda}\psi_{0}(\lambda)|\partial_{\lambda}\psi_{0}(\lambda)\rangle\\ -\langle\partial_{\lambda}\psi_{0}(\lambda)|\psi_{0}(\lambda)\rangle\langle\psi_{0}(\lambda)|\partial_{\lambda}\psi_{0}(\lambda)\rangle,$ (4)
which has a clear relation to the quantum metric tensor, i.e., the real part of the quantum geometric tensor [28, 7]. Using non-degenerate perturbation theory, we can also derive the spectral form,
$\chi^{H_{1}}_{F}(\lambda)$ $=\sum_{n\neq 0}\frac{|\langle\psi_{n}(\lambda)|H_{1}|\psi_{0}(\lambda)\rangle|^{2}}{[E_{n}(\lambda)-E_{0}(\lambda)]^{2}},$ (5)
where the superscript now indicates which term we treat as driving the perturbation. Finally, we can derive an integral representation [9] that is useful in deriving both scaling relations [44, 18] and in generalizing the Gibbs state for QMC [18, 17, 19].
To proceed [18], we define the imaginary time evolved operator $H_{1}(\tau)\equiv e^{\tau H(\lambda)}H_{1}e^{-\tau H(\lambda)}$ and the connected correlation function,
$G^{H_{1}}_{0}(\tau;\lambda)\equiv\theta(\tau)(\langle\psi_{0}(\lambda)|H_{1}(\tau)H_{1}|\psi_{0}(\lambda)\rangle\\ -\langle\psi_{0}(\lambda)|H_{1}|\psi_{0}(\lambda)\rangle^{2}),$ (6)
for $\theta(\tau)$ the Heaviside step function. Dropping $\lambda$ dependence for simplicity, inserting the resolution of the identity between $H_{1}$ and $e^{-\tau H}$ and simplifying, we find
$G^{H_{1}}_{0}(\tau;\lambda)=\theta(\tau)\sum_{n\neq 0}e^{\tau(E_{0}-E_{n})}|\langle\psi_{0}|H_{1}|\psi_{n}\rangle|^{2}.$ (7)
Taking the inverse Fourier transform,
$\widetilde{G}_{0}^{H_{1}}(\omega)\equiv\int_{\mathds{R}}e^{-i\tau\omega}G^{H_{1}}_{0}(\tau)=\sum_{n\neq 0}\frac{|\langle\psi_{0}|H_{1}|\psi_{n}\rangle|^{2}}{E_{n}-E_{0}+i\omega},$ (8)
we see a strong resemblance between this formula and Eq. 5. Namely, we recognize,
$\chi^{H_{1}}_{F}(\lambda)=i\frac{\mathrm{d}\widetilde{G}_{0}^{H_{1}}(\omega)}{\mathrm{d}\omega}=\int_{0}^{\infty}\tau G^{H_{1}}_{0}(\tau;\lambda)\mathrm{d}\tau,$ (9)
which is the sought after integral representation. This form relates $\chi_{F}$ to the dynamical response of the system to $H_{1}$, revealing its physical content [9, 44, 18].
A similar story holds for ES [18]. Applying non-degenerate perturbation theory, we find the second derivative of energy is given by
$\chi^{H_{1}}_{E}(\lambda)$ $=2\sum_{n\neq 0}\frac{|\langle\psi_{n}(\lambda)|H_{1}|\psi_{0}(\lambda)\rangle|^{2}}{E_{n}(\lambda)-E_{0}(\lambda)},$ (10)
which—aside from the pre-factor of $2$—differs from Eq. 5 only by the denominator. By inspection of Eq. 10 and Eq. 8,
$\chi^{H_{1}}_{E}(\lambda)=2\widetilde{G}^{H_{1}}_{0}(\omega=0)=2\int_{0}^{\infty}G^{H_{1}}_{0}(\tau;\lambda)\mathrm{d}\tau.$ (11)
With these integral expressions now derived, we can compare the ES and FS, Eqs. 11 and 9 more directly. Evidently, the important difference is that the FS involves a linear factor of $\tau$ in the integrand, which arises from the difference of the square in the denominators of Eqs. 10 and 5. Intuitively, one might expect that leads to the FS having a sharper peak at quantum critical points, and this holds rigorously by performing a standard scaling analysis [28, 47, 44]. Considering $L^{d}$ particles on a $d-$dimensional lattice that undergoes a second-order QPT at critical point $\lambda_{c},$ one can derive the scaling relations [18],
$L^{-d}\chi^{H_{1}}_{E,0}(\lambda)$ $\sim|\lambda-\lambda_{c}|^{-\alpha}\sim L^{(2/\nu)-(d+z)}$ (12)
$L^{-d}\chi^{H_{1}}_{F,0}(\lambda)$ $\sim|\lambda-\lambda_{c}|^{-(2-d\nu)}\sim L^{(2/\nu)-d}.$ (13)
Here, the correlation length diverges as $\xi\sim|\lambda-\lambda_{c}|^{-\nu},$ $z$ is the dynamical critical exponent, and $\alpha$ is defined through the hyperscaling formula $2-\alpha=\nu(d+z)$. Evidently, ES diverges whenever $\nu<2/(d+z)$ and the FS whenever $\nu<2/d$, and hence the FS is a stronger indicator of quantum criticality [18].
These groundstate concepts have also been successfully extended to mixed states [7, 8, 28], particularly thermal states [6, 18, 22]. A thermal state of a given Hamiltonian $H$ is given by
[…]
$G^{H_{1}}(\tau;\lambda;\beta)=\theta(\tau)(\langle H_{1}(\tau)H_{1}\rangle-\langle H_{1}\rangle^{2}),$ (15)
with which we can derive the integral forms [18],
$\chi^{H_{1}}_{E}(\lambda;\beta)$ $=\int_{0}^{\beta}G^{H_{1}}(\tau;\lambda;\beta)\mathrm{d}\tau$ (16)
$\chi^{H_{1}}_{F}(\lambda;\beta)$ $=\int_{0}^{\beta/2}\tau G^{H_{1}}(\tau;\lambda;\beta)\mathrm{d}\tau.$ (17)
These are particularly convenient for QMC [18, 17, 19]. The finite temperature connected correlator is periodic about $\beta/2$, so we can equivalently write
$\chi^{H_{1}}_{E}(\lambda;\beta)=2\int_{0}^{\beta/2}G^{H_{1}}(\tau;\lambda;\beta)\mathrm{d}\tau,$ (18)
for which it becomes apparent $\chi^{H_{1}}_{E}(\lambda;\beta)\rightarrow\chi^{H_{1}}_{E,0}(\lambda)$ and $\chi^{H_{1}}_{F}(\lambda;\beta)\rightarrow\chi^{H_{1}}_{F,0}(\lambda)$ as $\beta\rightarrow\infty$. In this sense, the finite temperature susceptibilities are a strict generalization of the groundstate susceptibilities, so we will henceforth write only $\chi_{E}$ and $\chi_{F}$ when our discussion does not rely on specific parameters at play, i.e. choice of driving term, $H_{1}$. A longer discussion connecting Eq. 17 to a more general notion of the Bures metric divergence at quantum critical points between mixed thermal states [6] is provided in Sec. II.B of Ref. [18].
III Numerical demonstrations
Before deriving formal estimators, we demonstrate our method through a series of numerical experiments using open source PMR-QMC code [32] that builds upon prior work [42, 23] (see Appendix A for basics of PMR-QMC). Each model is a spin-$1/2$ system defined in terms of the standard Pauli operators $X$, $Y$, and $Z$. We use the notation $X_{i}$, $Y_{i}$, and $Z_{i}$ to denote Pauli $X$, $Y$, and $Z$ operators acting on spin $i$, respectively.
We begin with illustrative experiments on well-known models, emphasizing agreement with expected results and, where applicable, reproducing prior plots [18, 19]. We then present a more challenging experiment involving an ensemble of random models, for which our approach still accurately estimates ES and FS. Notably, for this latter case, we are unaware of any other method capable of studying the model successfully. Yet, our single code can handle all of these models with only straightforward adjustments to input files [32].
III.1 A simple two spin model
We begin by studying,
$H(\lambda)=Z_{1}Z_{2}+\gamma(X_{1}+X_{2})+\lambda(Z_{1}+Z_{2}),$ (19)
first introduced and studied in a quantum sensing experiment on an NMR system via the fidelity approach (we choose $\gamma=0.1$[48]. By design, this model exhibits an avoided level crossing near the critical points $\lambda\approx\pm 1$, where the ES and FS are expected to peak. Our QMC approach correctly locates these peaks by computing the FS as shown in Fig. 1. In fact, as shown, our approach correctly matches direct numerical calculations for each temperature tested. Furthermore, our plot shows expected convergence of $\chi_{F}^{\beta}\rightarrow\chi_{F}$ as $\beta\rightarrow\infty$. At $\beta=20$ specifically, the QMC results closely match the exactly computed $T=0$ results. This verifies the basic veracity of our method.
Next, we remark that prior QMC approaches are generally incapable of estimating the FS, as they must typically restrict themselves to models for which $H_{1}\propto\text{diag}\left(H\right)$ or $H_{1}\propto H-\text{diag}\left(H\right)$ [17, 18, 19, 22]. Despite the simplicity of this model, it does not satisfy this condition. Yet, our method easily handles this model without even needing to write a specialized code for this case.
Figure 1: FS for the two spin model, Eq. (19), considered in Ref. [48] for different inverse temperatures, $\beta$. Data from QMC are points with 95% confidence intervals from bootstrapping over 5 independent random seeds, whereas direct numerical calculations are shown as lines. Naturally, there is no QMC data for $\beta=\infty$.
III.2 A square lattice TFIM
Secondly, we consider the well-known and extensively studied [18, 49, 50] transverse-field Ising model (TFIM),
$H_{\text{TFIM}}(h)=H_{J}+hH_{h}=-\sum_{\langle i,j\rangle}X_{i}X_{j}-h\sum_{i=1}^{n}Z_{i},$ (20)
on a square lattice with periodic boundary conditions (PBC). For our purposes, we focus our efforts on a simple $4\times 4$ model at fixed $\lambda=1$ and compute the ES and FS as a function of inverse temperature for either $H_{J}$ or $hH_{h}$ as the “driving term.” Notably, these results are actually not unconditional ES and FS as defined in Eqs. 16 and 17, but rather, the result of restricting computations only to the positive parity subspace of the TFIM. We refer interested readers to Appendix B for details on this methodological trick, but we remark that our approach is a non-trivial convenience of PMR-QMC that highlights its versatility.
Figure 2: We compute ES and FS restricted to the positive parity subspace (see App. B) of a $4\times 4$ square TFIM with PBC and $\lambda=1$ as a function of $\beta$. Our FS results are consistent with a similar computation performed with SSE-QMC [18].
The results are shown in Fig. 2 and demonstrate three things. Firstly, they showcase the ability of our method to consider different driving term choices as well as compute things in a fixed parity subspace. Secondly, they help verify the correctness of our approach, as the FS plot correctly matches results independently calculated with SSE-QMC [18]. Finally, they correctly reveal an expected feature of the ES and FS. Namely, one might naively expect that, for the purposes of detecting a quantum critical point, one is free to view either the static term $H_{J}$ or the driving term $hH_{h}$ in Eq. 20 as the perturbation.
In the $T=0$ limit, this is correct since
$\chi_{F,0}^{H_{J}}=h^{2}\chi_{F,0}^{hH_{h}},\ \ \ \chi_{E,0}^{H_{J}}=h^{2}\chi_{E,0}^{hH_{h}},$ (21)
follows from plugging in $H_{J}=H-hH_{h}$ into Eqs. 10 and 5 alongside the observation $\langle\psi_{n}|H|\psi_{0}\rangle=0$ for all $n\neq 0$. In these plots, we have chosen $h=1$, and correspondingly, we find that both the ES and FS converge for either perturbation choice ($H_{J}$ or $hH_{h})$ when $\beta$ is large enough. Yet for any finite $\beta$, the relation is less trivial, and in fact, one finds that there is no simple relation that holds across all inverse temperatures [18]. Methodologically, this provides a different way to test for empirical convergence of the finite temperature ES and FS to groundstate values.
So far, we have considered simple models of fixed size. To connect these ideas to the study of QPTs in the infinite system limit, we remark that one can simply compute the ES and FS for increasingly larger system sizes. This is done, for example, in Fig. 4 of Ref. [18]. As expected, this results in an increasingly sharper peak in both the ES and FS as $h$ approaches the known critical point $h_{c}\approx 3.044$ [50]. In addition, one can perform finite size scaling and data collapse in accordance with expected scaling relations, Eqs. 12 and 13, to extract the critical exponent $\nu\approx 0.6301$ [49, 18]. Expressed differently, our method is straightforwardly compatible with studying QPTs in the infinite size limit with enough computational effort, which we show more directly for the XXZ model next.
III.3 A square lattice XXZ model
Thirdly, we study the antiferromagnetic XXZ model on a square lattice with PBC,
$H=\frac{J_{z}}{4}\sum_{\langle i,j\rangle}Z_{i}Z_{j}+\frac{\lambda}{4}\sum_{\langle i,j\rangle}(X_{i}X_{j}+Y_{i}Y_{j}),$ (22)
which relates to a hard-core boson model [51, 52] via the Matsubara-Matsuda transformation [53]. The finite-temperature phase diagram of this model has been studied intensely [51], and various facts of this diagram have been replicated by FS approaches with exact diagonalization [54] and SSE-QMC [19]. From a QPT perspective, this model prefers Néel order in the XY plane when $\lambda$ dominates and an antiferromagnetic Ising groundstate when $J_{z}$ dominates. The point $\lambda=J_{z}$ is a known quantum critical point known as the Heisenberg point [19, 51].
Figure 3: We compute ES and FS for the XXZ model in Eq. 22 on a square lattice with PBC as a function of temperature, $T$, for different model sizes. Our FS results are consistent with a similar computation performed with SSE-QMC [19].
Furthermore [19, 51], it is known that thermal fluctuations at finite temperature destroy the antiferromagnetic Ising phase at a second-order phase transitions. Fixing $\lambda=1,J_{z}=1.5$ and varying temperature $T$, one can search for the expected critical point $(T/\lambda)_{c}\approx 0.75$ [19, 51]. Using the ES and FS, we search for this critical point in Fig. 3, and the empirical peak as we scale system size agrees with expectations. The FS plot in particular is consistent with results obtained in Ref. [19]. More broadly, these experiments show that our method can scale to large system sizes (which are necessary for finite size scaling in the study of QPTs) and also can study finite temperature phase transitions.
III.4 A random ensemble
Finally, we consider an ensemble of random 100-spin Hamiltonians generated by random unitary rotations of Eq. (19) (see Appendix C for full details of this family). By unitary invariance, observables for these random models agree with the simple two qubit model despite each $H_{U_{i}}=U_{i}H(\lambda)U_{i}^{\dagger}$ and associated driving term $U_{i}H_{1}U^{\dagger}_{i}$ being quite complicated, containing hundreds of random Pauli terms over 100 spins. Using standard QMC approaches, it is not possible to study this ensemble, but it is straightforward to try in PMR-QMC, and the results agree with known expectations as shown in Fig. 4.
(a) The imaginary time correlator, $G(\tau)$, has the expected symmetry around $\beta/2=10.$
(b) The ES correctly predicts critical points despite the $\beta=20$ curve not quite converging to the $\beta=\infty$ curve.
(c) The FS correctly predicts critical points more sharply than ES, and the $\beta=20$ curve matches the $\beta=\infty$ curve.
Figure 4: Comparison of QMC estimates with exact values for an ensemble of 100–spin Hamiltonians generated by random rotations of Eq. (19). Detailed information on the ensemble can be found in Appendix C. QMC estimates include 95% confidence intervals across 10 models, where each calculation employs a different random seed. In all cases, the QMC estimates show excellent agreement with exact finite $\beta$ curves.
Specifically, PMR-QMC estimators consistently show excellent agreement with exact numerical results for the imaginary-time correlator $G(\tau)$, the ES, and the FS, as shown in Fig. 4. Furthermore, our results confirm established characteristics of these observables; for instance, $G(\tau)$ exhibits symmetry about $\beta/2$ (see Fig. 1 in Ref. [18] and Fig. 7 in Ref. [19]).
IV Our estimators
Given a Hamiltonian of the form $H(\lambda)=H_{0}+\lambda H_{1}$, we can estimate the ES and FS via Eqs. 16 and 17 provided we can estimate $\langle H_{1}\rangle,\int_{0}^{\beta}\langle H_{1}(\tau)H_{1}\rangle$ and $\int_{0}^{\beta/2}\tau\langle H_{1}(\tau)H_{1}\rangle$. In permutation matrix representation QMC (PMR-QMC) [29, 23, 24, 25], we can derive estimators for each of these quantities independent of the details of $H$, and hence, for arbitrarily complicated driving terms $H_{1}$ [26]. This generality is what allowed us to, using a single estimator code, estimate ES and FS for the variety of models in Section III. In this section, we justify this claim by presenting explicit estimators and discussing their properties and computational complexity.
IV.1 PMR-QMC background and notation
A complete justification for our estimators requires substantial background in the permutation matrix representation (PMR) and PMR-QMC which we discuss in Appendix A (see also Refs. [26, 23]). The main derivations are also technical and lengthy and hence contained in Appendix D. Here, we provide a summary mainly to define the necessary notation and basic ideas.
[…]
$\mathcal{Z}=\sum_{z,S_{\bm{i}_{q}}}w_{(z,S_{\bm{i}_{q}})}\equiv\sum_{\mathcal{C}}w_{\mathcal{C}},$ (23)
for basis states $z$ and products of permutations $S_{\bm{i}_{q}}=P_{\bm{i}_{q}}\cdot\ldots\cdot P_{\bm{i}_{1}}$ for $q\geq 0$ and each $P_{\bm{i}_{k}}\in\{P_{j}\}$, and $w_{\mathcal{C}}$. Exploiting the fact that $G$ is Abelian and only those terms contribute where $S_{\bm{i}_{q}}$ evaluates to the identity, one can devise an automatic generation of a set of QMC update moves such that the resulting Markov chain is ergodic and satisfies detailed balance for essentially any physical Hamiltonian [23, 24, 25, 43]. This is the reason we are able to study a wide variety of models with a fixed code.
A PMR-QMC estimator $O_{\mathcal{C}}$ is a function of $\mathcal{C}$ such that
$\langle O\rangle=\frac{\sum_{\mathcal{C}}w_{\mathcal{C}}O_{\mathcal{C}}}{\sum_{\mathcal{C}}w_{\mathcal{C}}},$ (24)
which we denote as $O_{\mathcal{C}}\mathrel{\widehat{=}}\langle O\rangle$ to be read “$O_{\mathcal{C}}$ estimates $\langle O\rangle$” for convenience. Some of the present authors derived and explored the properties of a wide variety of estimators in PMR-QMC generally [26]. The basic idea is that since both $H$ and $O$ can always be decomposed in terms of the same group basis $G$, then one can always derive an expression for $O_{\mathcal{C}}$. In many cases of interest, such as if $O=H_{0}$ or $O=H_{1}$ for any bi-partition of $H$, then this formal expression leads to a faithful estimator in practical simulation [26]. This is why our estimators for the correlator, ES, and FS are independent of the choice of bi-partition.
To discuss our estimators more in-depth, we must introduce some standard PMR-QMC notation. First, we denote |$\ket{z_{k}}\equiv S_{\bm{i}_{k}}\ket{z}$ as the basis state resulting from applying the first $k$ permutations from $S_{\bm{i}_{q}}$. Next, we denote the diagonal energies $E_{z_{k}}\equiv\langle z_{k}|D_{0}|z_{k}\rangle.$ Writing $D_{\bm{i}_{k}}$ as the diagonal associated to $P_{\bm{i}_{k}}$, we can define the off-diagonal hopping term, $D_{(z,S_{\bm{i}_{q}})}\equiv\prod_{k}\langle z_{k}|D_{\bm{i}_{k}}|z_{k}\rangle$. Finally, we denote $e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}$ as the divided difference of the exponential $e^{-\beta x}$ with respect to the multiset of diagonal energies given [56, 57]. These thus defined, we can write the standard PMR-QMC result
$w_{(z,S_{\bm{i}_{q}})}\equiv D_{(z,S_{\bm{i}_{q}})}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]},$ (25)
as the efficiently computable PMR-QMC weight. We remark that the divided difference of the exponential (DDE) is especially important in our estimator derivations, and a detailed explanation of the DDE is given in Section A.3.
IV.2 Pure diagonal estimators
For simplicity, we first assume $H_{1}\propto D_{0}$, i.e., as in the TFIM studied in Section III.2. Using the properties of the PMR-QMC, we can readily derive the estimators,
$E_{z}$ =^$\mathrel{\widehat{=}}\langle D_{0}\rangle$ (26)
and
$\begin{split}\frac{E_{z}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}E_{z_{j}}e^{-\tau[E_{z_{j}},\ldots,E_{z_{q}}]}\\ \times\ e^{-(\beta-\tau)[E_{z_{0}},\ldots,E_{z_{j}}]}&\mathrel{\widehat{=}}\langle D_{0}(\tau)D_{0}\rangle\end{split}$ (27)
as we show in Appendix D. These estimators involve only sums and products of (1) diagonal energies $E_{z_{j}}$ and (2) divided differences of these diagonal energies—the same quantities that appear in the PMR-QMC weight, Eq. 25. As such, they are both readily computable in a PMR-QMC simulation in $O(1)$ and $O(q^{2})$ time, respectively (see Ref. [26] and Section A.4 to understand the time complexity). Utilizing both of these estimators, we can estimate $G^{D_{0}}(\tau;\lambda;\beta)$ via Eq. (15).
Furthermore, the only non-trivial $\tau$ dependence in these estimators is contained in the divided difference expressions of Eq. 27. Amazingly (see Lemma 1), we can readily derive the exact integral relation
$\int_{0}^{\beta}e^{-\tau[x_{j+1},\ldots,x_{q}]}e^{-(\beta-\tau)[x_{0},\ldots,x_{j}]}\mathrm{d}\tau=-e^{-\beta[x_{0},\ldots,x_{q}]}$ (28)
Combined with Eq. 27, we have thus derived
$\frac{-E_{z}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}E_{z_{j}}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}},E_{z_{j}}]}\\ \mathrel{\widehat{=}}\int_{0}^{\beta}\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau$ (29)
which gives an estimator for $\chi_{E}^{D_{0}}$ (see Eq. (16)) when combined with Eq. 26. As shown in Appendix D, a similar but more complicated derivation yields (see Lemma 3),
$\frac{E_{z}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}E_{z_{j}}\sum_{r=0}^{j}e^{-\frac{\beta}{2}[E_{z_{0}},\ldots,E_{z_{r}}]}\\ \times\sum_{m=j}^{q}e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q}},E_{z_{j}},E_{z_{m}}]}\mathrel{\widehat{=}}\int_{0}^{\beta/2}\tau\langle D_{0}(\tau)D_{0}\mathrm{d}\tau\rangle$ (30)
which gives an estimator for $\chi_{F}^{D_{0}}$ via Eq. (17). These integrated quantities and hence the ES and FS can be evaluated in $O(q^{2})$ and $O(q^{4})$ time, respectively.
Several comments are in order. Firstly, our derivations for ES and FS involve first deriving an estimator for the correlator. We then extend this estimator to the ES and FS by an analytical and exact evaluation of the divided difference integrals, i.e., no numerical integration is required. Secondly, these specific estimators are for $H_{1}\propto D_{0}$, but we can readily derive similar estimators for arbitrary $H_{1}$ (see Section IV.4). Even in the general case, the complexity to estimate the correlator, ES, and FS are $O(q^{2}),O(q^{2})$ and $O(q^{4})$, respectively. Thirdly, the complexity for the ES and FS can be improved in certain special cases as we discuss next.
IV.3 Improved ES and FS estimators
In the previous section, we derived $O(q^{2})$ and $O(q^{4})$ estimators for the ES and FS for $H_{1}\propto D_{0}$. These estimators can be improved to $O(1)$ and $O(q^{3})$ time in this case (as well as if $H_{1}\propto H-D_{0})$, which we state as theorems.
Theorem 1 (An $O(1)$ ES estimator).
The ES for $H_{1}\propto D_{0}$ or $H_{1}\propto H-D_{0}$ can be estimated in $O(1)$ time via
$\beta E_{z_{0}}^{2}+qE_{z_{0}}+\mathbf{1}_{q>0}\beta E_{z_{0}}\frac{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}]}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\\ \mathrel{\widehat{=}}\int_{0}^{\beta}\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau,$ (31)
for $\mathbf{1}_{q>0}$ zero if $q\leq 0$ and one if $q>0$.
Theorem 2 (An $O(q^{3})$ estimator for FS).
The FS for $H_{1}\propto D_{0}$ or $H_{1}\propto H-D_{0}$ can be computed in $O(q^{3})$ time via
$\mathbf{1}_{q=0}\frac{(\beta E_{z_{0}})^{2}}{8}+\mathbf{1}_{q>0}\frac{E_{z_{0}}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{r=0}^{q}e^{-\frac{\beta}{2}[E_{z_{0}},\ldots,E_{z_{r}}]}\\ \times\Bigg{(}\frac{\beta^{2}}{8}E_{z_{r}}e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q}}]}+\frac{\beta^{2}}{8}e^{-\frac{\beta}{2}[E_{z_{r+1}},\ldots,E_{z_{q}}]}-\\ \sum_{i=r+1}^{q}(i-r)e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q}},E_{z_{i}}]}\Bigg{)}\mathrel{\widehat{=}}\int_{0}^{\beta/2}\tau\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau,$ (32)
for $\mathbf{1}_{q>0}$ zero if $q\leq 0$ and one if $q>0$.
Our proof, given in complete detail in Appendix D, amounts to stating and proving several novel divided difference relations that can be used to simplify Eq. 29 and Eq. 30. While this naively only simplifies the $H_{1}\propto D_{0}$ case, the $H_{1}\propto H-D_{0}$ case follows by linearity and the derivation of a few $O(1)$ estimators for average energy $\langle H\rangle$ and similar quantities.
Stated abstractly, the constraint $H_{1}\propto D_{0}$ or $H_{1}\propto H-D_{0}$ may seem restrictive, but recall that this applies to the TFIM and XXZ models studied in Section III. More generally, many well studied models in condensed matter physics have such simple driving terms that are either purely diagonal or purely off-diagonal. As such, we have derived a constant time ES estimator for a wide variety of models.
IV.4 General case estimators
Generally, $H_{1}$ may contain both diagonal and off-diagonal terms, i.e., Eq. 19 or the random ensemble studied in Section III. This is not a fundamental issue for PMR-QMC, as whenever $H$ is in PMR form with $R$ term, $H=\sum_{j}D_{j}P_{j},$ we can always write $H_{1}=\sum_{j}\Lambda_{j}D_{j}P_{j}$ for $\Lambda_{j}$ diagonal matrices by the group structure of the PMR (see Appendix E.1). Observables of this type are said to be in canonical PMR form, and can be estimated correctly [26]. Namely, we can readily estimate $\langle H_{1}\rangle$ and $\langle H_{1}(\tau)H_{1}\rangle$ using expressions that generalize Eqs. 26 and 27.
For example, we can readily derive the estimator [26],
$\delta_{P_{l}}^{(q)}\frac{\Lambda_{l}(z)e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}]}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\mathrel{\widehat{=}}\langle\Lambda_{l}D_{l}P_{l}\rangle,$ (33)
from which we can estimate $\langle H_{1}\rangle$ by linearity. Here, $\delta_{P_{l}}^{(q)}$ equals $1$ if the $q^{\text{th}}$ or final permutation in $S_{\bm{i}_{q}}$ is $P_{l}$ and is 00 otherwise, and we have used the shorthand $\Lambda_{l}(z)\equiv\langle z|\Lambda_{l}|z\rangle.$ As for $\langle H_{1}(\tau)H_{1}\rangle$, we can also derive an estimator for $\langle(\Lambda_{k}D_{k}P_{k})(\tau)\Lambda_{l}D_{l}P_{l}\rangle$ that—up to small technicalities—contains the same $\tau$ dependence as Eq. 27 used to estimate $\langle D_{0}(\tau)D_{0}\rangle.$ Hence, even for arbitrary $H_{1}$, we can extend our estimator for $\langle H_{1}(\tau)H_{1}\rangle$ to estimators for $\chi_{E}^{H_{1}}$ and $\chi_{F}^{H_{1}}$ by our novel divided difference integral relations alone.
Specific expressions for these estimators and more discussion are provided in Appendix E for interested readers. The functional difference between these general estimators and those for $H_{1}\propto D_{0}$ or $H_{1}\propto H-D_{0}$ is that we estimate things term-by-term and exploit linearity. As such, the complexity estimation also depends on the number of non-trivial terms contained in $H_{1}$. Assuming there are $R$ non-trivial terms, the complexity of estimating the ES and FS in the most general case is then $O(R^{2}q^{2})$ and $O(R^{2}q^{4})$, respectively. In practice, the average complexity is much better, as we do not need to compute the full estimators whenever various $O(1)$ delta function conditions are not satisfied (see Eq. 33). Consequently, we were able to, for example, study the random ensemble with hundreds of terms successfully in Section III.
V Conclusions and future work
We have developed a universal black-box, finite-temperature quantum Monte Carlo (QMC) approach for studying quantum phase transitions (QPTs). Specifically, we derived PMR-QMC estimators for energy susceptibility (ES) and fidelity susceptibility (FS), which serve as general-purpose, order-parameter-free tools for locating quantum critical points. At the same time, given a Hamiltonian, we can compute a set of ergodic QMC updates that satisfy detailed balance.
Thanks to the generality of both the PMR-QMC framework and our PMR-based derivations, we successfully studied various aspects of QPTs, ES, and FS for a transverse-field Ising model, an XXZ model, and even a random ensemble—all using a single code. These demonstrations required no model-specific modifications to the source code—only a specification of the Hamiltonian and the desired partition, both of which can be arbitrary. As such, our method and publicly available open-source code are designed to be accessible and user friendly.
Of note, our method can estimate the ES in constant time for models with a simple driving term such as in the TFIM or XXZ model. Though the ES is expected to be a weaker indicator of criticality than the FS in general, we have found it to be sufficient in all models studied in this work. Hence, our approach is not only general but also very efficient for many systems of interest in condensed matter physics.
There are several promising avenues for future research. Foremost, we aim to apply our approach to physically relevant models with non-standard driving terms. We also intend to extend our method to bosonic, fermionic, and high-spin systems of current interest.
We hope the results of this work will provide a valuable and versatile tool for condensed matter physicists and researchers studying quantum phase transitions. Simulation code as well as data and analysis scripts are open source [32].
Acknowledgments
This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0063. All material, except scientific articles or papers published in scientific journals, must, in addition to any notices or disclaimers by the Contractor, also contain the following disclaimer: Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency (DARPA). N.E. was partially supported by the U.S. Department of Energy (DOE) Computational Science Graduate Fellowship under Award No. DE-SC0020347 and the ARO MURI grant W911NF-22-S-000 during parts of this work.
Appendix A Review of technical PMR-QMC background
A.1 Permutation matrix representation
The permutation matrix representation (PMR) [37] is a general matrix decomposition of square matrices [26] with several applications in condensed matter physics and quantum information [29, 37, 33, 24, 23, 58, 59, 60, 61, 25, 43]. Functionally, we say a square matrix $H$ is in PMR form provided it is written $H=D_{0}+\sum_{j}D_{j}P_{j}$ for each $D_{j}$ a diagonal matrix and $P_{j}$ permutations with no fixed points [37]. More rigorously [26], this is guaranteed provided $\{P_{j}\}\subset G$ for $G$ a permutation group with simply transitive group action on computational basis states, {|$\{\ket{z}\}$. The most important consequence of this form is that products of permutations $P_{i}P_{j}$ have no fixed points unless $P_{i}P_{j}=\mathds{1}$. A through definition of the PMR with several examples is contained in the work by some of the present authors [26].
A.2 Off-diagonal series expansion
[…]
$=\sum_{z}\sum_{n=0}^{\infty}\sum_{\{C_{\bm{i}_{n}}\}}\frac{f^{(n)}(0)}{n!}\langle z|C_{\bm{i}_{n}}|z\rangle,$ (35)
where in the second line we sum over all operator sequences consisting of $n$ products of $D_{0}$ and $D_{j}P_{j}$ terms, which we denote $\{C_{\bm{i}_{n}}\}$. The multi-index $\bm{i}_{n}\equiv(i_{1},\ldots,i_{n})$ denotes the ordered sequence, i.e. $\bm{i}_{3}=(3,0,1)$ indicates the sequence $C_{\bm{i}_{3}}=(D_{1}P_{1})D_{0}(D_{3}P_{3})$, read from right to left. More generally, each $i_{k}\in\{0,\ldots,M\}$ denotes a single term from the PMR form $H=\sum_{j=0}^{M}D_{j}P_{j}$.
For convenience, we can separate the contributions from diagonal operators $D_{j}$ from off-diagonal permutations, $P_{j}$ which yields [31, 30] the following complicated expression,
[…]
which is justified in prior works [31, 30]. We now briefly summarize the notation, which will be used throughout. Firstly, $S_{\bm{i}_{q}}=P_{\bm{i}_{q}}\cdot\ldots\cdot P_{\bm{i}_{1}}$ denotes a product over $q$ permutations, each taken from $\{P_{j}\}_{j=1}^{M}$, i.e., the multiset indices are now $\bm{i}_{j}\in\{1,\ldots,M\}$. Next, we denote |$\ket{z_{0}}\equiv\ket{z}$ and |$\ket{z_{k}}\equiv P_{\bm{i}_{k}}\cdot\ldots\cdot P_{\bm{i}_{1}}\ket{z}$. This allows us to define the “diagonal energies” as $E_{z_{k}}\equiv\langle z_{k}|H|z_{k}\rangle=\langle z_{k}|D_{0}|z_{k}\rangle$ (recall |$\ket{z_{k}}$ is a basis for $D_{0}$ not of $H$) and the off-diagonal “hopping strengths” $D_{(z,S_{\bm{i}_{q}})}\equiv\prod_{k=1}^{q}\langle z_{k}|D_{\bm{i}_{k}}|z_{k}\rangle$. The sum over $\sum_{\sum k_{i}}$ is again the set of weak partitions of $n-q$ into $q+1$ integers.
This complicated expression can be greatly simplified in two ways. Firstly, by the no-fixed points property of the PMR [26], $\langle z|S_{\bm{i}_{q}}|z\rangle=1$ if and only if $S_{\bm{i}_{q}}=\mathds{1}$ and is zero otherwise. Secondly, as first noticed in earlier works [31, 30], the diagonal term in parenthesis can be identified with the so-called divided difference of $f$ [56, 57]. Together, we can write
[…]
A.3 Divided differences
We briefly review the technical details of the divided difference, inspired by the treatment in Refs. [31, 30, 26]. The divided difference of any holomorphic function $f(x)$ can be defined over the multiset $[x_{0},\ldots,x_{q}]$ using a contour integral [57, 56],
$f[x_{0},\ldots,x_{q}]\equiv\frac{1}{2\pi i}\oint\limits_{\Gamma}\frac{f(x)}{\prod_{i=0}^{q}(x-x_{i})}\mathrm{d}x,$ (38)
for $\Gamma$ a positively oriented contour enclosing all the $x_{i}$’s. Several elementary properties utilized in PMR-QMC [31, 29] follow directly from this integral representation, or by invoking Cauchy’s residue theorem. For example, $f[x_{0},\ldots,x_{q}]$ is invariant to permutations of arguments, the definition reduces to Taylor expansion weights when arguments are repeated,
$f[x_{0},\ldots,x_{q}]=f^{(q)}(x)/q!,\ \ \ x_{0}=x_{1}\ldots=x_{q}=x,$ (39)
and whenever each $x_{i}$ is distinct, we find
$f[x_{0},\ldots,x_{q}]=\sum_{i=0}^{q}\frac{f(x_{i})}{\prod_{k\neq i}(x_{i}-x_{k})},$ (40)
the starting definition in Refs. [31, 30, 29]. Each of these divided difference definitions can be shown to satisfy the Leibniz rule [56],
$(f\cdot g)[x_{0},\ldots,x_{q}]=\sum_{j=0}^{q}f[x_{0},\ldots,x_{j}]g[x_{j},\ldots,x_{q}]=\sum_{j=0}^{q}g[x_{0},\ldots,x_{j}]f[x_{j},\ldots,x_{q}]$ (41)
which is particularly important in our derivations.
Yet another useful way to view the divided difference for our work is to derive its power series expansion,
$f[x_{0},\ldots,x_{q}]=\sum_{m=0}^{\infty}\frac{f^{(q+m)}(0)}{(q+m)!}\sum_{\sum k_{j}=m}\prod_{j=0}^{q}x_{j}^{k_{j}},$ (42)
where the notation $\sum_{\sum k_{j}=m}$ is a shorthand introduced in Refs. [30, 31] which represents a sum over all weak integer partitions of $m$ into $q+1$ parts. More explicitly, it is an enumeration over all vectors $\bm{k}$ in the set $\{\bm{k}=(k_{0},\ldots,k_{q}):k_{i}\in\mathds{N}_{0},\sum_{j=0}^{q}k_{j}=m\}$, where $\mathds{N}_{0}$ is the natural numbers including zero. To derive the series expansion, we first Taylor expand $f(x)$ inside the contour integral,
$f[x_{0},\ldots,x_{q}]$ $=\sum_{n=0}^{\infty}\frac{f^{(n)}(0)}{n!}\left(\frac{1}{2\pi i}\oint_{\Gamma}\frac{x^{n}}{\prod_{i=0}^{q}(x-x_{i})}\right)=\sum_{n=0}^{\infty}\frac{f^{(n)}(0)}{n!}[x_{0},\ldots,x_{q}]^{n},$ (43)
where we have introduced the shorthand $[x_{0},\ldots,x_{q}]^{k}\equiv p_{k}[x_{0},\ldots,x_{q}]$ for $p_{k}(x)=x^{k}$ as introduced in Refs. [30, 31]. The divided difference of a polynomial has a closed form expression most easily written with the change of variables $n\rightarrow q+m$,
$[x_{0},\ldots,x_{q}]^{q+m}=\begin{cases}0&m<0\\ 1&m=0\\ \sum_{\sum_{k_{j}}=m}\prod_{j=0}^{q}x_{j}^{k_{j}}&m>0,\end{cases}$ (44)
as noted in the same notation in Refs. [30, 31] and derived with a different notation in Ref. [56]. Together, these two observations yield Eq. 42.
As with previous PMR works [31, 29], we are also especially interested in the divided difference of the exponential (DDE) where we utilize the shorthand notation $e^{t[x_{0},\ldots,x_{q}]}\equiv f[x_{0},\ldots,x_{q}]$ for $f(x)=e^{tx}$. Replacing $f^{(q+m)}(0)$ with $t^{q+m}$ gives the power series expansion of the DDE, as derived in the appendices of Refs. [31, 30]. Replacing the variable $x\rightarrow\alpha x$ in Eq. 38, we find the rescaling relation,
$\alpha^{q}e^{t[\alpha x_{0},\ldots,\alpha x_{q}]}=e^{\alpha t[x_{0},\ldots,x_{q}]},$ (45)
which is useful in numerical schemes [55, 62] and in computing the Laplace transform. In particular, combining Eq. 45 with Eq. (11) of Ref. [63] with $P_{m}(x)=1$, we find
$\mathcal{L}\left\{e^{\alpha t[x_{0},\ldots,x_{q}]}\right\}=\frac{\alpha^{q}}{\prod_{j=0}^{q}(s-\alpha x_{j})},$ (46)
where $\mathcal{L}$ denotes the Laplace transform from $t\rightarrow s$. One can alternatively derive Eq. 46 by directly performing the integration to the contour integral definition in Eq. 38, e.g. by Taylor expanding $e^{\alpha tx}$ and re-summing term-by-term, legitimate by the uniform convergence of the exponential.
A.4 Computational complexity of estimators in PMR-QMC
Throughout this work, we refer to complexity of evaluating PMR-QMC estimators. For interested readers, more details and examples of tracking complexity in PMR-QMC estimation are contained in Sec. VIII of Ref. [26]. Here, we give a brief summary of how time is accounted.
[…]
Appendix B Parity restricted PMR-QMC for the TFIM
In our transverse-field Ising model (TFIM) numerical experiments of Section III.2, we computed the ES and FS restricted to the positive parity subspace of the TFIM. We briefly discuss the details of our methodology. First, we study a rotated TFIM for which the transverse field term is $\sum_{i}Z_{i}$, Eq. 20. In doing so, the parity operator,
$\mathcal{P}=\prod_{i=1}^{N}Z_{i}$ (47)
commutes with the TFIM, $[H,\mathcal{P}]=0$. Hence, the parity $P=\pm 1$ is a good quantum number, and it is well known that the groundstate of the TFIM lies in the $P=+1$ parity sector by the Perron-Frobenius theorem [18].
For our rotated TFIM choice, $\mathcal{P}$ is purely diagonal. For both SSE-QMC [18] and PMR-QMC (see Eq. 23), computational basis elements |$\ket{z}$ are part of the QMC configuration. At any given time in simulation, then, one can query the eigenvalue of $\mathcal{P}\ket{z}$, which is the instantaneous parity $P_{\mathcal{C}}$. Note that, in fact, |$\ket{z}$ is a bitstring, so the parity can readily be obtained by determining if there are an odd number of $1$’s in |$\ket{z}$ or not. In the SSE-QMC approach [18], the authors simply keep track of $P_{\mathcal{C}}$ during simulation and post-select on measurement outcomes.
In contrast, we perform PMR-QMC sampling restricted to the positive parity subspace the entire simulation. This is possible because the set of PMR-QMC updates is standardized in terms of the permutation group that comprises $H$. A careful look at the updates for spin-1/2 systems [23] reveals that there are only two updates that can change a configurations parity: a classical update or a block swap. In our TFIM implementation [32], we simply choose an initial |$\ket{z}$ with positive parity and only allow classical updates or block swaps that preserve parity. Both can be altered to preserve parity without post-selection, which is therefore more efficient than the SSE approach.
In a standard classical update, a random bit is flipped. In our code, we simply flip two random bits instead, which preserves parity. In a standard block swap, the classical state is also updated. Specifically, the permutation string $S_{\bm{i}_{q}}$ is split into two $S_{\bm{i}_{q}}=S_{2}S_{1}$ for $S_{1}=P_{i_{k}}\cdots P_{i_{1}}$ and $S_{2}=P_{i_{q}}\cdots P_{i_{k+1}}$ for $k\in\{1,\ldots,q-1\}$ chosen randomly. The classical state at position $k$ is then |$\ket{z^{\prime}}=S_{1}\ket{z}$ with energy $E_{z^{\prime}}$ The updated block-swap configuration is then $\mathcal{C^{\prime}}=\{\ket{z^{\prime}},S_{1}S_{2}\}$, which swaps $S_{1}$ an $S_{2}$ while updating |$\ket{z}\rightarrow\ket{z^{\prime}}$. The parity |$\ket{z^{\prime}}$ can change based on $S_{1}$.
For the TFIM, however, each $P_{i_{l}}$ is just a 2-body $X$ string, i.e., $X_{1}X_{2}$, $S_{\bm{i}_{q}}=\mathds{1}$ if and only if each permutation appears an even number of times. Together, this means $S_{1}$ flips the parity of |$\ket{z^{\prime}}$ if and only if $k$ is odd. Hence, parity can be preserved by only sampling $k\in\{1,\ldots,q-1\}$ that is also even. This can easily be accomplished, and is what we do in our code [32]. Together, these small changes ensure we stay in the $P=+1$ parity sector during the entire PMR-QMC simulation.
Appendix C Additional details on generating special unitaries
Consider $U=\sum_{k=1}^{L}c_{k}Q_{k}$ for $\bm{c}\in\mathds{R}^{L}$ and each $Q_{k}$ a Pauli matrix over $n$ spins (we use $Q$ to distinguish from the permutation $P_{j}$ in the PMR form of $H$). By construction, we can efficiently form the matrix $UH(\lambda)U^{\dagger}$ for $H(\lambda)=Z_{1}Z_{2}+0.1(X_{1}+X_{2})+\lambda(Z_{1}+Z_{2}),$ our prototypical example Hamiltonian from Ref. [48], whenever $L$ is polynomial in $n$. This is because for each term in $H$, we need only perform $L^{2}$ Pauli conjugations, i.e, compute terms like $Q_{k}X_{1}Q_{l}$, which are themselves computable in $O(n)$ time using a look-up table or a symplectic inner product technique as in Clifford simulations.
In addition, $U$ can easily be shown to be unitary whenever $||\bm{c}||_{2}=1$ and $\{Q_{k}\}_{k=1}^{L}$ all anti-commute (see Eq. (6) in [65]). The largest maximal set of anti-commuting Paulis is $12n+$ 1, and there is an $O(n)$ algorithm to generate a canonical set of this type [66]. Let $\mathcal{Q}$ denote this canonical set of $201$ Paulis for a 100–spin system, which, in particular, includes $X^{\otimes 100}$. We can now describe our algorithm to generate 10 random unitaries such that $H_{U_{i}}\equiv U_{i}H(\lambda)U_{i}^{\dagger}$ satisfies the four properties described in the main text.
To begin with, we generate 50 random unitaries by (a) selecting $L\in[1,\ldots,201]$ uniformly at random, (b) construct a random list of anti-commuting Paulis of the form $\{X^{\otimes 100}\}\cup\{Q_{i}\}_{i=2}^{L}$ for each $Q_{i}$ drawn uniformly without replacement from $\mathcal{Q}$, (c) generating $c_{i}\sim\mathcal{N}(0,1)$ for $i=1,\ldots,L$ independently and then renormalizing. In step (b), we always include $X^{\otimes 100}$ to ensure each $H_{U_{i}}(\lambda)\equiv U_{i}H(\lambda)U^{\dagger}_{i}$ has non-trivial support on all 100 spins. By construction, $H_{U_{i}}$ has hundreds of Paulis terms, and yet, it is easy to generate and store. What remains to show is that we can find a subset of 10 without a severe sign problem.
To our knowledge, there is no simple calculation to check for the sign problem for these models a priori. So in practice, we simply ran a PMR-QMC simulation for each model for $\beta=5$ and $\lambda=1,$ which approximately the critical point for this model. The PMR-QMC code in Ref. [42, 23] automatically tracks the average and variance of the sign of the QMC weight, so we then simply post-selected those models with the highest average sign. Empirically, we find the worst average sign in this set of 10 is $0.9983\pm 0.0006$ (this is $2\sigma$), so this subset is essentially sign problem free.
Appendix D Derivations of PMR-QMC estimators for pure diagonal or pure off-diagonal driving
We derive the ES and FS estimators for the $H_{1}\propto D_{0}$ case stated in Sections IV.2 and IV.3. Most notably, we prove Theorems 1 and 2. We then substantiate the claim that these same estimators also straightforwardly carry over to the $H_{1}\propto H-D_{0}$ case.
D.1 Derivation strategy
Recall from the main text (or see Ref. [26]) that a PMR-QMC estimator $O_{\mathcal{C}}$ is a function of $\mathcal{C}$ such that
$\langle O\rangle=\frac{\sum_{\mathcal{C}}w_{\mathcal{C}}O_{\mathcal{C}}}{\sum_{\mathcal{C}}w_{\mathcal{C}}},$ (48)
for $w_{\mathcal{C}}=D_{(z,S_{\bm{i}_{q}})}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}$ the generalized PMR-QMC Boltzmann weight. For convenience, we write $O_{\mathcal{C}}\mathrel{\widehat{=}}\langle O\rangle$ to be read “$O_{\mathcal{C}}$ estimates $\langle O\rangle$.” Our first goal is to derive an estimator for the ES when $H_{1}\propto D_{0}$.
In most practical applications, $H_{1}\propto D_{0}$ follows from the relation $\lambda H_{1}=D_{0}$, i.e., the driving term multiplied by $\lambda$ is the diagonal portion of $H$. As such, $\langle H_{1}\rangle=\langle D_{0}\rangle/\lambda$, and more generally, it is enough to simply estimate $\langle D_{0}\rangle$ and related quantities. To this end, we observe
$\chi_{E}^{D_{0}}(\lambda;\beta)=\int_{0}^{\beta}(\langle D_{0}(\tau)D_{0}\rangle-\langle D_{0}\rangle^{2})\mathrm{d}\tau.$ (49)
gives the ES (see Eq. 16) up to a multiplicative correction, i.e., $\chi_{E}^{H_{1}}=\chi_{E}^{D_{0}}/\lambda^{2}$ (or more generally just $\chi_{E}^{H_{1}}\propto\chi_{E}^{D_{0}})$. Thus, we focus on attention on estimating $\chi_{E}^{D_{0}}$.
Our strategy is to first derive estimators for $\langle D_{0}\rangle$ and $\langle D_{0}(\tau)D_{0}\rangle$ in Section D.2. As we shall see, $E_{z}\mathrel{\widehat{=}}\langle D_{0}\rangle$ does not depend on $\tau$. Then, in Section D.3, we show $\int_{0}^{\beta}\langle D_{0}(\tau)D_{0}\rangle$ can be evaluated exactly and analytically into a valid estimator via novel divided difference relations. Together, these results give an $O(q^{2})$ estimator for ES. In Section D.3, we further simplify our estimator into a constant time $O(1)$ ES estimator using additional novel divided difference relations. This completes most of the work of proving Theorem 1.
By a similar chain of reasoning,
$\chi_{F}^{D_{0}}(\lambda;\beta)=\int_{0}^{\beta/2}\tau(\langle D_{0}(\tau)D_{0}\rangle-\langle D_{0}\rangle^{2})\mathrm{d}\tau,$ (50)
gives the FS (see Eq. 17) up to a multiplicative correction, and our strategy remains similar. Namely, we show in Section D.5 that $\int_{0}^{\beta/2}\tau\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau$ can also be evaluated exactly and analytically via novel divided difference relations. This results in an $O(q^{4})$ estimator for FS. By yet additional novel divided difference relations, we reduce this to $O(q^{3})$ operations in Section D.6. This l lcompletes most of the work of proving Theorem 2.
Finally, we show how all the estimators we derived for the $H\propto D_{0}$ case carry over straightforwardly to the $H_{1}\propto H-D_{0}$ case and with the same complexities in Section D.7 thus completing the justification for Theorems 1 and 2.
D.2 Estimators for $\langle D_{0}\rangle$ and $\langle D_{0}(\tau)D_{0}\rangle$
[…]
$E_{z}\mathrel{\widehat{=}}\langle D_{0}\rangle,$ (51)
is an $O(1)$ estimator (see Section A.4).
Next, we derive an estimator for $\langle D_{0}(\tau)D_{0}\rangle$. One approach is by direct off-diagonal series expansion and simplification [26]. Here, we argue that direct Leibniz rule logic (Eq. 41), we can correctly conclude,
$\langle z|D_{0}e^{-\tau H}D_{0}e^{-(\beta-\tau)H}|z\rangle=E_{z}\sum_{j=0}^{q}E_{z_{j}}e^{-\tau[E_{z_{j}},\ldots,E_{z_{q}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,E_{z_{j}}]}.$ (52)
The basic idea is simple: without the middle $D_{0}$ term, we can safely apply the Leibniz rule to the off-diagonal expansion since $e^{-\tau H}$ and $e^{-(\beta-\tau)H}$ commute. Accounting for the non-commuting $D_{0}$, however, is simple—we just pick up a factor of $E_{z_{j}}$ from the $D_{0}\ket{z_{j}}$ as shown. Hence, we find
$(D_{0}(\tau)D_{0})_{\mathcal{C}}\equiv\\ \frac{E_{z}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}E_{z_{j}}e^{-\tau[E_{z_{j}},\ldots,E_{z_{q}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,E_{z_{j}}]}\mathrel{\widehat{=}}\langle D_{0}(\tau)D_{0}\rangle.$ (53)
as claimed in Eq. 27.
D.3 Energy susceptibility in quadratic time
We shall show
$\frac{-E_{z_{0}}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}E_{z_{j}}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}},E_{z_{j}}]}\mathrel{\widehat{=}}\int_{0}^{\beta}\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau.$ (54)
Combined with Eqs. 51 and 49, this gives an estimator for $\chi_{E}^{D_{0}}$ that requires $O(q^{2})$ operations (see Section A.4). By inspection of Eq. 53, showing Eq. 54 amounts to proving the following integral relation.
Lemma 1 (Convolution lemma).
$\int_{0}^{\beta}e^{-\tau[x_{j+1},\ldots,x_{q}]}e^{-(\beta-\tau)[x_{0},\ldots,x_{j}]}\mathrm{d}\tau=-e^{-\beta[x_{0},\ldots,x_{q}]}.$ (55)
Proof.
For convenience, we define the functions
$f(t)$ $=e^{-t[x_{j+1},\ldots,x_{q}]}$ (56)
$g(t)$ $=e^{-t[x_{0},\ldots,x_{j}]}$ (57)
The convolution of these functions,
$(f*g)(t)$ $=\int_{0}^{t}f(\tau)g(t-\tau)\mathrm{d}\tau$ (58)
is by construction the integral we want to evaluate for $t=\beta$. Let $\mathcal{L}\left\{f(t)\right\}$ denote the Laplace transform of $f(t)$ from $t\rightarrow s$. By the convolution property of the Laplace transform and Eq. 46, we find
$\mathcal{L}\left\{(f*g)(t)\right\}$ =$=\mathcal{L}\left\{f(t)\right\}\mathcal{L}\left\{g(t)\right\}$ (59)
$=\left(\frac{(-1)^{q-j-1}}{\prod_{l=j+1}^{q}(s+x_{l})}\right)\left(\frac{(-1)^{j}}{\prod_{m=0}^{j}(s+x_{m})}\right)=\frac{(-1)^{q-1}}{\prod_{l=0}^{q}(s+x_{l})}=\mathcal{L}\left\{-e^{-t[x_{0},\ldots,x_{q}]}\right\}.$ (60)
Taking the inverse Laplace transform of the first and final expression proves the claimed integral relation. ∎
Inspection of Eq. 53 and Eq. 55 shows the validity of Eq. 54 immediately. As an aside, we remark that a direct proof by series expanding both DDE via Eq. 42, integrating term-by-term, regrouping, and simplifying is also possible.
D.4 Energy susceptibility in constant time
The ES estimator in Eq. 54 requires $O(q^{2})$ operations to evaluate, but it can be simplified into an $O(1)$ estimator and validate Theorem 1 in a few steps. We begin with a simple proposition.
Proposition 1 (Re-scaling relation).
$e^{-\tau[x_{0},\ldots,x_{q}]}=(-\tau)^{q}e^{[-\tau x_{0},\ldots,-\tau x_{q}]}.$ (61)
Proof.
By the change of variables $y=-\tau z$ in Eq. (38), we find the claimed result. ∎
We can now simplify the sum in Eq. 54 into a compact derivative form.
Lemma 2 (A weighted, repeated argument sum simplification).
$\sum_{j=0}^{q}x_{j}e^{-\tau[x_{0},\dots,x_{q},x_{j}]}=\left(\tau\frac{\partial}{\partial\tau}-q\right)e^{-\tau[x_{0},\dots,x_{q}]}.$ (62)
Proof.
We prove this by induction.
Base case. When $q=0$, we have $x_{0}e^{-\tau[x_{0},x_{0}]}=x_{0}\partial e^{-\tau x_{0}}/\partial x_{0}=-\tau x_{0}e^{-\tau x_{0}}$, so Eq. (62) holds.
Induction step. By Prop. 1, we write the right-hand side as,
$(\tau\partial_{\tau}-(q+1))(-\tau)^{q+1}e^{[-\tau x_{0},\ldots,-\tau x_{q+1}]}$ $=-(-\tau)^{q+2}\partialderivative{\tau}e^{[-\tau x_{0},\ldots,-\tau x_{q+1}]}$ (63)
$=-(-\tau)^{q+2}\frac{1}{2\pi i}\oint_{\Gamma}\partialderivative{\tau}\frac{e^{z}}{\prod_{i=0}^{q+1}(z+\tau x_{i})}\mathrm{d}z=\sum_{j=0}^{q+1}x_{j}e^{-\tau[x_{0},\ldots,x_{q+1},x_{j}]}$ (64)
Therefore, Eq. (62) is true for all $q$. ∎
To turn this simplification into a useful result for PMR-QMC estimation, we employ the following proposition.
Proposition 2 (Parametric derivative).
$\partialderivative{\tau}e^{-\tau[x_{0},\ldots,x_{q}]}=-x_{0}e^{-\tau[x_{0},\ldots,x_{q}]}-e^{-\tau[x_{1},\ldots,x_{q}]}.$ (65)
Proof.
First, define $g(x)=-xe^{-\tau x}$ for convenience. By Eq. (38),
$\partialderivative{\tau}e^{-\tau[x_{0},\ldots,x_{q}]}\equiv\partialderivative{\tau}\frac{1}{2\pi i}\oint_{\Gamma}\frac{e^{-\tau z}}{\prod_{i=0}^{q}(z-x_{i})}\mathrm{d}z=\frac{1}{2\pi i}\oint_{\Gamma}\frac{-ze^{-\tau z}}{\prod_{i=0}^{q}(z-x_{i})}\mathrm{d}z\equiv g[x_{0},\ldots,x_{q}].$ (66)
Employing the Leibniz rule, Eq. 41, we get the desired result. ∎
Corollary 1 (Computing the weighted, repeated argument sum).
$\sum_{j=0}^{q}x_{j}e^{-\tau[x_{0},\dots,x_{q},x_{j}]}=\left\{\begin{array}[]{ll}(-x_{0}\tau-q)e^{-\tau[x_{0},\dots,x_{q}]}-\tau e^{-\tau[x_{1},\dots,x_{q}]},&\text{\rm for}\quad q>0,\\ -\tau x_{0}e^{-\tau x_{0}},&\text{\rm for}\quad q=0.\end{array}\right.$ (67)
Proof.
This follows from applying Prop. 2 to Lem. 2. ∎
We now prove the main part of Theorem 1—that we have an $O(1)$ estimator for ES—since
$\beta E_{z_{0}}^{2}+qE_{z_{0}}+\mathbf{1}_{q>0}\beta E_{z_{0}}\frac{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}]}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\mathrel{\widehat{=}}\int_{0}^{\beta}\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau.$ (68)
Proof.
The original ES estimator in Eq. (54) can be simplified using Eq. (67). The first two terms follow by simple algebraic manipulation. The final term follows by algebra and periodicity, i.e., $E_{z_{0}}=E_{z_{q}}$ for a valid set of PMR-QMC diagonal energies. After a PMR-QMC update, we have $O(1)$ access to $\beta$, each $E_{z_{j}},$ and each successive divided difference $e^{-\beta E_{z_{0}}},e^{-\beta[E_{z_{0}},E_{z_{1}}]},\ldots,e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}$, which includes $e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}]}$. ∎
D.5 Fidelity susceptibility in quartic time
We shall now show
$\frac{E_{z_{0}}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}E_{z_{j}}\sum_{r=0}^{j}e^{-\frac{\beta}{2}[E_{z_{0}},\ldots,E_{z_{r}}]}\sum_{m=j}^{q}e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q}},E_{z_{j}},E_{z_{m}}]}\mathrel{\widehat{=}}\int_{0}^{\beta/2}\tau\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau,$ (69)
which gives rise to an $O(q^{4})$ estimator (see Section A.4) for $\chi_{F}^{D_{0}}$ in Eq. 50 when combined with Eq. 51. As with the ES, this essentially reduces to proving divided difference relations since we need only integrate the expression in Eq. 53. To this end, we first state and prove a helpful proposition.
Proposition 3 (Repeated argument sum simplification).
$\sum_{j=0}^{q}e^{-\tau[x_{0},\ldots,x_{q},x_{j}]}=-\tau e^{-\tau[x_{0},\ldots,x_{q}]}$ (70)
Proof.
Denote $f(\tau)\equiv e^{-\tau[x_{0},\ldots,x_{q}]}$ and recall $\mathcal{L}\left\{e^{-\tau[x_{0},\ldots,x_{q}]}\right\}=(-1)^{q}/(\prod_{i=0}^{q}(s+x_{i})).$ By the derivative property of the Laplace transform,
$\mathcal{L}\left\{-\tau f(\tau)\right\}=\partialderivative{s}\mathcal{L}\left\{f(\tau)\right\}=\sum_{j=0}^{q}\frac{(-1)^{q+1}}{(s+x_{j})\prod_{i=0}^{q}(s+x_{i})}=\mathcal{L}\left\{\sum_{j=0}^{q}e^{-\tau[x_{0},\ldots,x_{q},x_{j}]}\right\},$ (71)
and the result follows by taking the inverse Laplace transform of the left-most and right-most expressions. ∎
Using this proposition and results we have shown previously, we can now prove the following integral relation.
Lemma 3 (Fidelity susceptibility integral).
$I\equiv\int_{0}^{\beta/2}\tau e^{-\tau[x_{j+1},\ldots,x_{q}]}e^{-(\beta-\tau)[x_{0},\ldots,x_{j}]}\mathrm{d}\tau=\sum_{r=0}^{j}e^{-\frac{\beta}{2}[x_{0},\ldots,x_{r}]}\sum_{m=j+1}^{q}e^{-\frac{\beta}{2}[x_{r},\ldots,x_{q},x_{m}]},$ (72)
Proof.
This relation follows from the Leibniz rule (Eq. 41),  Proposition 3, and  Eq. 55 applied in order,
$I$ $=\sum_{r=0}^{j}e^{-(\beta/2-\tau)[x_{0},\ldots,x_{r}]}\int_{0}^{\beta/2}\tau e^{-\tau[x_{j+1},\ldots,x_{q}]}e^{-(\beta/2-\tau)[x_{r},\ldots,x_{j}]}\mathrm{d}\tau$ (73)
$=\sum_{r=0}^{j}\sum_{m=j+1}^{q}e^{-(\beta/2-\tau)[x_{0},\ldots,x_{r}]}\int_{0}^{\beta/2}e^{-\tau[x_{j+1},\ldots,x_{q},x_{m}]}e^{-(\beta/2-\tau)[x_{r},\ldots,x_{j}]}\mathrm{d}\tau$ (74)
$=\sum_{r=0}^{j}e^{-(\beta/2-\tau)[x_{0},\ldots,x_{r}]}\sum_{m=j+1}^{q}e^{-\frac{\beta}{2}[x_{r},\ldots,x_{q},x_{m}]},$ (75)
where in the final line we use $\beta\rightarrow\beta/2$ when employing Eq. 55. ∎
This lemma is already sufficient to demonstrate Eq. 69, but to make things clearer, we state the following corollary first.
Corollary 2.
Lemma 3 is general for any multiset $[x_{0},\ldots,x_{q}]$ partitioned at arbitrary index $j$. As a specific special case of interest now, we consider $[x_{j+1},\ldots,x_{q}]$ replaced with $[x_{j},\ldots,x_{q}]$,
$I^{\prime}\equiv\int_{0}^{\beta/2}\tau e^{-\tau[x_{j},\ldots,x_{q}]}e^{-(\beta-\tau)[x_{0},\ldots,x_{j}]}\mathrm{d}\tau=\sum_{r=0}^{j}e^{-\frac{\beta}{2}[x_{0},\ldots,x_{r}]}\sum_{m=j}^{q}e^{-\frac{\beta}{2}[x_{r},\ldots,x_{q},x_{j},x_{m}]}.$ (76)
Proof.
This result follows from Lemma 3 with the main change being that the $m$ index now starts at $j$ instead of $j+1$. To explain the notation change, it is useful to simply repeat the above derivation, which proceeds as,
$I^{\prime}$ $=\sum_{r=0}^{j}e^{-(\beta/2-\tau)[x_{0},\ldots,x_{r}]}\int_{0}^{\beta/2}\tau e^{-\tau[x_{j},\ldots,x_{q}]}e^{-(\beta/2-\tau)[x_{r},\ldots,x_{j}]}\mathrm{d}\tau$ (77)
$=\sum_{r=0}^{j}\sum_{m=j}^{q}e^{-(\beta/2-\tau)[x_{0},\ldots,x_{r}]}\int_{0}^{\beta/2}e^{-\tau[x_{j},\ldots,x_{q},x_{m}]}e^{-(\beta/2-\tau)[x_{r},\ldots,x_{j}]}\mathrm{d}\tau$ (78)
$=\sum_{r=0}^{j}e^{-(\beta/2-\tau)[x_{0},\ldots,x_{r}]}\sum_{m=j+1}^{q}e^{-\frac{\beta}{2}[x_{r},\ldots,x_{q},x_{j},x_{m}]},$ (79)
where the last multiset now contains an explicit extra $x_{j}$ from the concatenation of $[x_{j},\ldots,x_{q},x_{m}]$ and $[x_{r},\ldots,x_{j}]$ which both contain an explicit $x_{j}$. ∎
Inspection of Eqs. 53 and 76 shows the validity of Eq. 69 immediately.
D.6 Fidelity susceptibility in cubic time
The FS estimator in Eq. 69 can be evaluated in $O(q^{4})$ operations, but we can improve it to $O(q^{3})$ and validate Theorem 2 in a few steps. We begin with a some helpful propositions and lemmas, where the first step is to re-order the sum.
Proposition 4 (Re-ordering sums in FS estimator).
$\frac{E_{z_{0}}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{r=0}^{q}e^{-\frac{\beta}{2}[E_{z_{0}},\ldots,E_{z_{r}}]}\sum_{\begin{subarray}{c}j,m=r\\ j\leq m\end{subarray}}^{q}E_{z_{j}}e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q}},E_{z_{j}},E_{z_{m}}]}\mathrel{\widehat{=}}\int_{0}^{\beta/2}\tau\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau$ (80)
With the re-ordering, we now observe that the innermost sum can be simplified.
Lemma 4 (A weighted, repeated argument, double, less-than sum simplification).
$\sum_{\begin{subarray}{c}i,j=0\\ i\leq j\end{subarray}}^{q}x_{i}e^{-\tau[x_{0},\dots,x_{q},x_{i},x_{j}]}=\left(-\frac{\tau^{2}}{2}\frac{\partial}{\partial\tau}-\sum_{i=0}^{q}i\cdot\frac{\partial}{\partial x_{i}}\right)e^{-\tau[x_{0},\dots,x_{q}]}.$ (81)
Proof.
We prove this by induction.
Base case. When $q=0$, we have
$x_{0}e^{-\tau[x_{0},x_{0},x_{0}]}=x_{0}\cdot\frac{1}{2}\frac{\partial^{2}}{\partial x_{0}^{2}}e^{-\tau x_{0}}=\frac{\tau^{2}x_{0}}{2}e^{-\tau x_{0}}=-\frac{\tau^{2}}{2}\frac{\partial}{\partial\tau}e^{-\tau x_{0}},$
so Eq. (81) holds.
Induction step. It follows from the recursive definition of the divided differences that
$\sum_{\begin{subarray}{c}i,j=0\\ i\leq j\end{subarray}}^{q+1}x_{i}e^{-\tau[x_{0},\dots,x_{q+1},x_{i},x_{j}]}=\frac{1}{x_{0}-x_{1}}\sum_{\begin{subarray}{c}i,j=0\\ i\leq j\end{subarray}}^{q+1}x_{j}\left(e^{-\tau[x_{0},x_{2},\dots,x_{q+1},x_{i},x_{j}]}-e^{-\tau[x_{1},x_{2},\dots,x_{q+1},x_{i},x_{j}]}\right)=\\ =\frac{1}{x_{0}-x_{1}}\left(\left(\left(-\frac{\tau^{2}}{2}\frac{\partial}{\partial\tau}-\sum_{i=1}^{q}i\cdot\frac{\partial}{\partial x_{i+1}}\right)e^{-\tau[x_{0},x_{2},\dots,x_{q+1}]}+x_{0}e^{-\tau[x_{0},x_{2},\dots,x_{q+1},x_{0},x_{1}]}+x_{1}\sum_{j=1}^{q+1}e^{-\tau[x_{0},\dots,x_{q+1},x_{j}]}\right)-\right.\\ \left.-\left(\left(-\frac{\tau^{2}}{2}\frac{\partial}{\partial\tau}-\sum_{i=1}^{q}i\cdot\frac{\partial}{\partial x_{i+1}}\right)e^{-\tau[x_{1},x_{2},\dots,x_{q+1}]}+x_{0}\sum_{j=0}^{q+1}e^{-\tau[x_{0},x_{1},\dots,x_{q+1},x_{j}]}\right)\right)=\\ =\left(-\frac{\tau^{2}}{2}\frac{\partial}{\partial\tau}-\sum_{i=0}^{q}i\cdot\frac{\partial}{\partial x_{i+1}}\right)e^{-\tau[x_{0},\dots,x_{q+1}]}-\sum_{j=1}^{q+1}e^{-\tau[x_{0},\dots,x_{q+1},x_{j}]}=\\ =\left(-\frac{\tau^{2}}{2}\frac{\partial}{\partial\tau}-\sum_{i=0}^{q+1}i\cdot\frac{\partial}{\partial x_{i}}\right)e^{-\tau[x_{0},\dots,x_{q+1}]}.$ (82)
Therefore, Eq. (81) is true for all $q$. ∎
As with the ES simplification, this more compact expression involving a derivative can be converted into a valid PMR-QMC estimator using Proposition 2.
Corollary 3.
$\sum_{\begin{subarray}{c}i,j=0\\ i\leq j\end{subarray}}^{q}x_{i}e^{-\tau[x_{0},\dots,x_{q},x_{i},x_{j}]}=\left\{\begin{array}[]{ll}\left(\tau^{2}x_{0}/2\right)e^{-\tau[x_{0},\dots,x_{q}]}+(\tau^{2}/2)e^{-\tau[x_{1},\dots,x_{q}]}-\sum_{i=1}^{q}i\cdot e^{-\tau[x_{0},\dots,x_{q},x_{i}]},&\text{\rm for}\quad q>0,\\ (\tau^{2}x_{0}/2)e^{-\tau x_{0}},&\text{\rm for}\quad q=0.\end{array}\right.$ (83)
Proof.
This simplification follows from applying Prop. 2 to Eq. (81). ∎
We now prove the main part of Theorem 2—that we have an $O(q^{3})$ estimator for FS—since
$\mathbf{1}_{q=0}\frac{(\beta E_{z_{0}})^{2}}{8}+\mathbf{1}_{q>0}\frac{E_{z_{0}}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{r=0}^{q}e^{-\frac{\beta}{2}[E_{z_{0}},\ldots,E_{z_{r}}]}\Bigg{(}\frac{\beta^{2}}{8}E_{z_{r}}e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q}}]}+\frac{\beta^{2}}{8}e^{-\frac{\beta}{2}[E_{z_{r+1}},\ldots,E_{z_{q}}]}-\\ \sum_{i=r+1}^{q}(i-r)e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q}},E_{z_{i}}]}\Bigg{)}\mathrel{\widehat{=}}\int_{0}^{\beta/2}\tau\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau$ (84)
Proof.
The estimator follows from applying Eq. (83) to the estimator form given in Eq. (80), taking special care to correctly translate the sum from $i,j=0$ to $q$ on arguments $[x_{0},\ldots,x_{q}]$ to the sum of $j,m=r$ to $q$ on arguments $[E_{z_{r}},\ldots,E_{z_{q}}]$. The estimate $O(q^{3})$ comes from the observation that, for each $r$, we must do $O(q^{2})$ work. ∎
D.7 Off-diagonal case follows from diagonal case
We now consider the case $H_{1}\propto H-D_{0}$ and show it reduces to using the $H_{1}\propto D_{0}$ estimators. Let
$\Gamma\equiv H-D_{0},$ (85)
and define the associated ES and FS
$\chi_{E}^{\Gamma}(\lambda;\beta)$ $=\int_{0}^{\beta}(\langle\Gamma(\tau)\Gamma\rangle-\langle\Gamma\rangle^{2})\mathrm{d}\tau$ (86)
$\chi_{F}^{\Gamma}(\lambda;\beta)$ $=\int_{0}^{\beta/2}\tau(\langle\Gamma(\tau)\Gamma\rangle-\langle\Gamma\rangle^{2})\mathrm{d}\tau.$ (87)
As with the diagonal case, we find $\chi_{E}^{H_{1}}\propto\chi_{E}^{\Gamma}$ and $\chi_{F}^{H_{1}}\propto\chi_{F}^{\Gamma}$. Since $H=D_{0}+\sum_{j>0}D_{j}P_{j}$, then it follows that $\Gamma=\sum_{j>0}D_{j}P_{j}$. One approach to estimate the ES and FS in this case is to derive estimators for $\langle D_{j}P_{j}\rangle$ and $\langle(D_{l}P_{l})(\tau)D_{j}P_{j}\rangle$ and use linearity. While this does work, if $\Gamma$ contains $R$ terms, then it incurs an overhead of doing $R$ and $R^{2}$ estimations. This overhead can be avoided by using simple estimators for $H$ and $D_{0}$ quantities and linearity.
Firstly, $\langle\Gamma\rangle=\langle H-D_{0}\rangle=\langle H\rangle-\langle D_{0}\rangle$ by linearity. From prior work (see Eq.63 in Ref. [26]), we know
$E_{z_{q}}+\mathbf{1}_{q\geq 1}\frac{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}}]}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}=\begin{cases}E_{z_{q}}&q=0\\ E_{z_{q}}+\frac{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}}]}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}&q>0,\end{cases}\mathrel{\widehat{=}}\langle H\rangle$ (88)
where $\mathbf{1}_{q\geq 1}$ is the indicator function that is 0 when $q<1$ and 1 when $q\geq 1$. Since $E_{z}\mathrel{\widehat{=}}\langle D_{0}\rangle$, then by linearity we find,
$\mathbf{1}_{q\geq 1}\frac{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}}]}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\mathrel{\widehat{=}}\langle\Gamma\rangle.$ (89)
This estimator can be used to estimate $\langle\Gamma\rangle^{2}.$
Next, we observe by linearity and cyclicity of the trace,
$\langle\Gamma(\tau)\Gamma\rangle$ $=\langle H^{2}\rangle+\langle D_{0}(\tau)D_{0}\rangle-2\langle D_{0}H\rangle,$ (90)
which has the full non-trivial $\tau$ dependence in the familiar diagonal correlator. As such, we can estimate $\langle\Gamma(\tau)\Gamma\rangle$ and ES and FS integrals thereof provided we can derive estimators for $\langle H^{2}\rangle$ and $\langle D_{0}H\rangle$. By off-diagonal expansion manipulations, we know (e.g., Eq. 65 of Ref. [26]),
$E_{z_{q}}^{2}+\frac{\mathbf{1}_{q\geq 1}(E_{z_{q}}+E_{z_{q-1}})e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}]}+\mathbf{1}_{q\geq 2}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-2}}]}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\mathrel{\widehat{=}}\langle H^{2}\rangle$ (91)
and (see Eq. 70 in Ref. [26]),
$E_{z_{0}}\left(E_{z_{0}}+\mathbf{1}_{q\geq 1}\frac{e^{-\beta[E_{z_{1}},\ldots,E_{z_{q}}}]}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\right)\mathrel{\widehat{=}}\langle D_{0}H\rangle.$ (92)
Each of the static estimators, i.e., for $\langle D_{0}\rangle,\langle H\rangle,\langle H^{2}\rangle,and\langle D_{0}H\rangle$, can be computed in constant time $O(1)$. Integrals of these $\tau$-independent quantities are also trivial. Hence, the entire complexity in evaluating $\langle\Gamma(\tau)\Gamma\rangle$ is $O(q^{2})$ inherited from the complexity of computing $\langle D_{0}(\tau)D_{0}\rangle$. Similarly, the complexity of evaluating ES and FS is thus $O(1)$ and $O(q^{3})$ via the estimators we derived in Appendix D. Namely, we have now finished our proof of Theorems 1 and 2 by showing that the non-trivial parts of the ES and FS estimators for both the $H_{1}\propto D_{0}$ and $H_{1}\propto H-D_{0}$ are the same and hence have the same complexity.
Appendix E Most general driving term estimators
In the most general case, $H_{1}$ contains both diagonal and off-diagonal terms and does not satisfy $H_{1}\propto D_{0}$ or $H_{1}\propto H-D_{0}$. Nevertheless, we can characterize the general form of $H_{1}$ for arbitrary Hamiltonians and use that to derive estimators for ES and FS in this arbitrary case.
E.1 The PMR form of $H_{1}$
In the study of quantum phase transitions, we assume $H(\lambda)=H_{0}+\lambda H_{1}$ for some $[H_{0},H_{1}]\neq 0$. Regardless of the specifics, this Hamiltonian can always be in PMR form, $H(\lambda)=\sum_{j}D_{j}(\lambda)P_{j}$, where we choose $P_{0}=\mathbb{1}$ by convention. Here, we show that one can always write
$H_{1}=\sum_{j}\Lambda_{j}(\lambda)D_{j}(\lambda)P_{j}$ (93)
where $\Lambda_{j}(\lambda)$ are diagonal matrices and the $D_{j}(\lambda)P_{j}$ terms are directly taken from the PMR form of $H$. The consequence of this form is that there are no fundamental obstacles to estimating $\langle H_{1}\rangle,\langle H_{1}(\tau)H_{1}\rangle,$ and integrals thereof using PMR-QMC, and in fact, estimators have a relatively simple form. A full understanding of why this is true is subtle and entrenched in PMR-QMC details, and hence, out of the scope of this work. For additional context, we point to Section V.B in Ref. [23] and Ref. [26], which discuss measurements of arbitrary operators $O$ and what properties $O$ can have in relation to $H$ that prevent accurate estimation by a single PMR-QMC simulation. To reiterate our main point, proving Eq. (93) is one way to show that $H_{1}$ does not have any of these subtle problems that prevent estimation.
To begin with, we build up the PMR form of $H(\lambda)$ by first writing PMR forms for $H_{0}$ and $H_{1}$ separately. For simplicity, denote $P\in H_{0}$ to mean there exists a term $DP$ in the PMR expansion of $H_{0}$ where $D\neq 0$. With this notation, we can write,
$H_{0}=\sum_{P\in H_{0}}A_{P}P,\ H_{1}=\sum_{P\in H_{1}}B_{P}P\implies H(\lambda)=\sum_{P\in H_{0},P\in H_{1}}(A_{P}+\lambda B_{P})P+\sum_{P\in H_{0},P\notin H_{1}}A_{P}P+\sum_{P\notin H_{0},P\in H_{1}}\lambda B_{P}P.$ (94)
When compared with $H(\lambda)=\sum_{j}D_{j}(\lambda)P_{j}$, we can readily identify $\{P_{j}\}_{j}=\{P\in H_{0}\}\cup\{P\in H_{1}\}$ and $D_{j}$ as the diagonal in front of $P_{j}$, i.e., either $A_{j},\lambda B_{j},$or $(A_{j}+\lambda B_{j})$. Before showing Eq. (93) generally, we remark that in the extremely common cases that $H_{1}$ is either purely diagonal or purely off-diagonal, things simplify and Eq. (93) is obvious. In the pure diagonal case, $H_{1}=\frac{\mathbb{1}}{\lambda}D_{0}$, and in the pure off-diagonal case, $H_{1}=\sum_{j>0}\frac{\mathbb{1}}{\lambda}D_{j}$.
Now for the general case, consider $P\in H_{0},P\notin H_{1}$ where we have $D_{j}=A_{j}$ and choose $\Lambda_{j}=0$. For $P\notin H_{0},P\in H_{1}$, we have $D_{j}=B_{j}$ and choose $\Lambda_{j}=\mathbb{1}$. Finally, for $P\in H_{0},P\in H_{1}$, we have $D_{j}=(A_{j}+\lambda B_{j})$ and choose $\Lambda_{j}=(A_{j}+\lambda B_{j})^{+}B_{j}$, where $X^{+}$ denotes the pseudo-inverse of $X$, which prevents diving by zero. Before justifying the pseudo-inverse—which is more subtle than any discussion so far—we remark that this case is moot if $H_{1}$ is purely diagonal or purely off-diagonal, since such partitions have no $P$ that is contained in both $H_{0}$ and in $H_{1}$. To see why the pseudo-inverse works, suppose $\langle z|A_{j}+\lambda B_{j}|z\rangle=0$ which implies either $\langle z|A_{j}|z\rangle=\langle z|B_{j}|z\rangle=0$ or $A_{j}=-\lambda B_{j}$. In the first case, $\langle z|\Lambda_{j}B_{j}|z\rangle=0$ for any choice of $\Lambda_{j}$, so the pseudo-inverse works. The second-case is not relevant since it requires fine-tuning $\lambda$, and such exceptional points can be avoided in practice while still generating $\chi_{F}^{\beta}(\lambda)$ curves.
E.2 Estimating $\langle H_{1}\rangle$ and $\langle H_{1}(\tau)H_{1}\rangle$
In the prior section, we showed that we can write $H_{1}=\sum_{j}\Lambda_{j}D_{j}P_{j}$ for $\Lambda_{j}$’s diagonal matrices whenever $H=\sum_{j}D_{j}P_{j}$ and $P_{0}=\mathds{1}$. By linearity, it is sufficient to derive estimators for each $\langle\Lambda_{j}D_{j}P_{j}\rangle$ term, for which it is useful to separate diagonal term, $\Lambda_{0}$ from the general off-diagonal terms $\Lambda_{l}D_{l}P_{l}$ for $l>0$. In both cases, some of the present authors have derived the estimators (see Eqs.56 and 82) [26],
$\Lambda_{0}(z)$ =^$\mathrel{\widehat{=}}\langle\Lambda_{0}\rangle$ (95)
$\delta_{P_{l}}^{(q)}\frac{\Lambda_{l}(z)e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}]}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}$ =^$\mathrel{\widehat{=}}\langle\Lambda_{l}D_{l}P_{l}\rangle,$ (96)
and hence, we can evaluate $\langle H_{1}\rangle$ in $O(R)$ calls to these $O(1)$ estimators. Here, $\Lambda_{l}(z)\equiv\langle z|\Lambda_{l}|z\rangle$ and $\delta_{P_{l}}^{(q)}=1$ if the $q^{\text{th}}$ permutation in $S_{\bm{i}_{q}}$ is $P_{l}$ and is 00 otherwise. (Recall that products of permutations $S_{\bm{i}_{q}}$ partially make up PMR-QMC configurations as explained in Appendix A.)
Similarly, we can estimate $\langle H_{1}(\tau)H_{1}\rangle$ by evaluating all $R^{2}$ estimators of the form $\langle A(\tau)B\rangle$ for $A,B\in\{\Lambda_{0},\Lambda_{1}D_{1}P_{1},\ldots,\Lambda_{R-1}D_{R-1}P_{R-1}\}.$ Just as the above estimator, it is best to consider the four possible ways of combining a pure diagonal contribution with a pure off-diagonal one,
$\frac{\Lambda_{0}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}\Lambda_{0}(z_{j})e^{-\tau[E_{z_{j}},\ldots,E_{z_{q}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,E_{z_{j}}]}$ =^$\mathrel{\widehat{=}}\langle\Lambda_{0}(\tau)\Lambda_{0}\rangle$ (97)
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}\Lambda_{0}(z_{j})e^{-\tau[E_{z_{j}},\ldots,E_{z_{q-1}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,E_{z_{j}}]}$ =^$\mathrel{\widehat{=}}\langle(\Lambda_{k}D_{k}P_{k})(\tau)\Lambda_{0}\rangle$ (98)
$\frac{\Lambda_{0}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}\delta^{(j)}_{P_{l}}\Lambda_{l}(z_{j})e^{-\tau[E_{z_{j}},\ldots,E_{z_{q}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,E_{z_{j-1}}]}$ =^$\mathrel{\widehat{=}}\langle\Lambda_{0}(\tau)(\Lambda_{l}D_{l}P_{l})\rangle$ (99)
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}{\delta_{P_{l}}^{(j)}}{\Lambda_{l}(z_{j})}e^{-\tau[E_{z_{j}},\ldots,{E_{z_{q-1}}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,{E_{z_{j-1}}}]}$ =^$\mathrel{\widehat{=}}\langle(\Lambda_{k}D_{k}P_{k})(\tau)(\Lambda_{l}D_{l}P_{l})\rangle,$ (100)
where each estimator follows as different special cases of the derivation of Eq. 119 of Ref. [26]. As with the $\langle D_{0}(\tau)D_{0}\rangle$ estimator, each of these estimators requires $O(q^{2})$ operations, so the total complexity to evaluate $\langle H_{1}(\tau)H_{1}\rangle$ is simply $O(R^{2}q^{2})$.
Although these estimators are more complicated than the pure diagonal counterparts Section D.2, the big picture for ES and FS is the same. Namely, the only non-trivial $\tau$ dependence is contained in the divided differences of the correlator estimator, for which we can apply the divided difference integral relations we have already derived.
E.3 General ES estimator
As in the derivation of the diagonal ES estimator, we need only consider deriving an estimator for $\int_{0}^{\beta}\langle H_{1}(\tau)H_{1}\rangle\mathrm{d}\tau$ and combine it with our $\langle H_{1}\rangle$ estimators. Although $\langle H_{1}(\tau)H_{1}\rangle$ really consists of four different possible estimators, each of these contains an expression of the form $e^{-\tau[\ldots]}e^{-(\beta-\tau)[\ldots]}$ for which we can apply the convolution lemma, Lemma 1. For example, applying this lemma to Eqs. 97 and 100, we find
$\frac{\Lambda_{0}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}\Lambda_{0}(z_{j})e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}},E_{z_{j}}]}$ =^$\mathrel{\widehat{=}}\int_{0}^{\beta}\langle\Lambda_{0}(\tau)\Lambda_{0}\rangle\mathrm{d}\tau$ (101)
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}{\delta_{P_{l}}^{(j)}}{\Lambda_{l}(z_{j})}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}]}$ =^$\mathrel{\widehat{=}}\int_{0}^{\beta}\langle(\Lambda_{k}D_{k}P_{k})(\tau)(\Lambda_{l}D_{l}P_{l})\rangle\mathrm{d}\tau.$ (102)
The former clearly resembles $\langle D_{0}(\tau)D_{0}\rangle$ and is computable in $O(q^{2})$ time. Unfortunately, unlike the diagonal case simplification in Theorem 1, there does not appear any way to simplify the sum into an $O(1)$ expression since $\Lambda_{0}(z_{j})$ has no relation to $E_{z_{j}}$. On the other hand, the second estimator can be simplified. Namely, since the inner divided difference has no direct $j$ dependence,
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}{\delta_{P_{l}}^{(j)}}{\Lambda_{l}(z_{j})}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}]}=\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}]}\sum_{j=1}^{q-1}{\delta_{P_{l}}^{(j)}}{\Lambda_{l}(z_{j})},$ (103)
then this is actually an $O(q)$ expression. In total, this means evaluating $\chi_{E}^{H_{1}}$ for completely general $H_{1}$ takes $O(R^{2}q^{2})$ time. In the special case that $H_{1}$ has no pure diagonal term, however, our ES estimator needs $O(R^{2}q)$ effort instead.
E.4 General FS estimator
For the FS, the same logic as in the previous section applies except we employ the more complicated integral lemma, Lemma 3, to evaluate $\int_{0}^{\beta/2}\tau\langle H_{1}(\tau)H_{1}\rangle\mathrm{d}\tau$. Doing so to Eqs. 97 and 100, we find
$\frac{\Lambda_{0}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}\Lambda_{0}(z_{j})\sum_{r=0}^{j}e^{-\frac{\beta}{2}[E_{z_{0}},\ldots,E_{z_{j}}]}\sum_{m=j}^{q}e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q}},E_{z_{j}},E_{z_{m}}]}$ =^$\mathrel{\widehat{=}}\int_{0}^{\beta/2}\tau\langle\Lambda_{0}(\tau)\Lambda_{0}\rangle\mathrm{d}\tau$ (104)
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}{\delta_{P_{l}}^{(j)}}{\Lambda_{l}(z_{j})}\sum_{r=0}^{j-1}e^{-\frac{\beta}{2}[E_{z_{0}},\ldots,E_{z_{j}}]}\sum_{m=j-1}^{q-1}e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q-1}},E_{z_{m}}]}$ =^$\mathrel{\widehat{=}}\int_{0}^{\beta/2}\tau\langle(\Lambda_{k}D_{k}P_{k})(\tau)(\Lambda_{l}D_{l}P_{l})\rangle\mathrm{d}\tau,$ (105)
where neither can be simplified by our prior formulas or clear new insights. As such, both require $O(q^{4})$ effort to compute, and the overall complexity of a general FS estimator requires $O(R^{2}q^{4})$ time.